ri File Edit View History Bookmarks Window Help hu 12:04 AM 6. O -13 points My Notes Ask Your Te If a solution containing 0.16 M co32 Bro I-, and N s treated with Agt, in what order will the anions precipitate? (order your answer from first to last. Use the appropriate or symbol to separate substances in the list.) chem Pad Help Supporting Materials Periodic Table Physical Constants Supplemental Materials
Answer #1

Ksp of Ag2CO3 =8.5 ×10^-12

Ksp of AgBrO3 = 5.5 × 10^-5

Ksp of AgI = 8.5 × 10^-17

Ksp of AgN3 = 2 × 10^-8

Which having low Ksp that will precipitate first

Therefore the order is

AgI > Ag2CO3 > AgN3 > AgBrO3